If A is a square matrix then,
(A) A + AT is symmetric (B) AAT is skew - symmetric (C) AT + A is skew-symmetric(D) ATA is skew symmetric

Grade:

2 Answers

Arun
3 years ago
Dear student
 
let C = A + A^T
 
Now C^T = A^T + A 
 
hence C is a symmetric matrix.
 
So option A is correct.
 
Thanks and regards
Vikas TU
3 years ago
The determinant of a skew symmetric matrix of odd order is zero, which means skew symmetric matrix of an odd number are singular. Hence their inverse does not exist.